# Universal Hitchin moduli spaces

### v1

@@ 1. Introduction
The moduli space $\mathcal{R}$ of representations of the fundamental group of a closed oriented surface $\Sigma$ has an interesting interplay with classical Teichmüller theory. For the case of a compact Lie group $K$, a choice of complex structure $J$ on $\Sigma$ determines an homeomorphism of the moduli space $\mathcal{R}(K)$ of $K$-representations with a moduli space of holomorphic principal $G$-bundles over the compact Riemann surface $X_{J}=(\Sigma,J)$, via the Narasimhan–Seshadri theorem [NS] and its extension to principal bundles by Ramanathan [R1, R2]. Here, $G$ denotes the complex reductive Lie group given by the complexification of $K$. This way, the character variety $\mathcal{R}(K)$ determines a holomorphic fibration over the Teichmüller space $\mathcal{T}$ of $\Sigma$, given by complex structures on $\Sigma$ modulo the action of diffeomorphisms isotopic to the identity.
@@
(1.1) $\begin{array}[]{l}F_{A}-[\varphi,\tau(\varphi)]=0,\\ \bar{\partial}_{J,A}\varphi=0,\end{array}$
and subsequent work by Simpson [Si3, Si4]. Here $A$ is a connection on a principal $K$-bundle over $\Sigma$ and $\varphi\in\Omega^{1,0}(X_{J},E_{K}(\mathfrak{g}))$ is the Higgs field. The corresponding gauge-theoretical moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ has three different incarnations, as a character variety $\mathcal{R}_{c}(G)$ for a complex reductive Lie group $G$, as a moduli space space $\mathcal{M}^{\operatorname{Flat}}(G)$ of reductive flat $G$-connections, and as a moduli space of polystable $G$-Higgs bundles $\mathcal{M}(G)$. These three viewpoints on the moduli space of Hitchin’s equations determine a hyperkähler structure on $\mathcal{M}^{\operatorname{Hit}}(G)$. Interestingly, only the moduli space $\mathcal{M}(G)$ depends on the choice of a complex structure on $\Sigma$. Furthermore, the twistor space $Z\to\mathbb{CP}^{1}$ for the hyperkähler structure has generic complex structure $\mathcal{M}^{\operatorname{Flat}}(G)$, while the special points over $\mathbb{CP}^{1}$ correspond to $\mathcal{M}(G)$ and its conjugate.
In this paper we investigate metric aspects of the universal moduli space of solutions of Hitchin’s equations varying over the Teichmüller space $\mathcal{T}$. Our approach is gauge-theoretical in nature, and builds on the classical theory of Kähler fibrations [GLS, M] and in the symplectic interpretation of constant scalar curvature Kähler metrics [Do2, F]. We fix a symplectic structure $\omega$ on $\Sigma$ and consider the coupled system of equations
@@
where $J$ is a complex structure on $\Sigma$ and $S_{g}$ is the scalar curvature of the metric $g=\omega(,J)$. We denote by $\mathcal{U}^{Hit}(G)$ the moduli space of solutions of (1.2) modulo unitary gauge (see Section 5.4). Our first main result can be stated as follows (cf. Theorem 5.18 and Theorem 5.22):
@@
(1.3) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mu_{\curly{H}}^{-1}(0)/\curly{H}$
@@
(1.5) $\begin{split}\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}_{2}+\tfrac{\alpha}{2}\psi(J\dot{J}_{2})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right),\end{split}$
where $B$ is a positive-definite invariant metric on $\mathfrak{k}=\operatorname{Lie}K$. Hence, it is not obvious a priori that $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$. The key step is to undertake a gauge fixing for solutions of the coupled Hitchin’s equations (1.2), whereby the complex structure $\mathbb{I}$ and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space. Difficulties will arise, due to the fact that $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is neither a definite pairing nor non-degenerate.
@@
(1.7) $\mathcal{U}^{*}_{\alpha,\varepsilon}\to\mathcal{U}^{\operatorname{Flat}}(G)$
@@
and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [F, Section 4]).
Similarly as for the universal moduli space of Hitchin’s equations (1.2), the (possibly degenerate) moduli space Kähler form and metric admit explicit formulae upon suitable gauge fixing
(1.8) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v,v)&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2}))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1}))\wedge J(\psi(\dot{J}_{2}))).\end{split}$
From the previous expression, we observe that, at least formally, the restriction of $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ to the fibres over $\mathcal{T}$ coincides up to scaling with Hitchin’s symplectic structure $\omega_{\mathbf{J}}$. A complete proof of this requires a better understanding of the particular gauge fixing mechanism which we use to construct the moduli space complex structure $\mathbb{J}$ (see Proposition 4.9), and we leave it as an open question.
@@
We are far from even grasping the solution of this proposal, and we leave it as an open question for future studies.
Motivation for the present work comes from the programme initiated by the first three authors in [AGG] more than 15 years ago based on the 2009 PhD Thesis of the second author [GF]. In that paper, we consider a pair $(X,E)$ consisting of a holomorphic principal $G$-bundle $E$ over a compact complex manifold $X$, where $G$ is a complex reductive group, and study certain coupled equations for a Kähler metric on $X$ and a reduction of structure group of $E$ to a maximal compact subgroup $K\subset G$. These equations, that we named Kähler–Yang–Mills equations, appear naturally as moment map equations for the action of the extended gauge group of the $K$-bundle $E_{K}$ over the underlaying smooth manifold to $X$ equipped with a symplectic form $\omega$. The extension of the gauge group of $E_{K}$ is given in this case by the group of Hamiltonian symplectomorphisms defined by $\omega$. This theory is in some sense a combination of the usual Yang–Mills theory and the Donaldson–Fujiki theory on Kähler manifolds, interpreting the constant scalar curvature as the moment map for the action of the group of Hamiltonian symplectomorphisms on the space of complex structures compatible with the symplectic structure $\omega$.
@@
In early joint work with the fourth author, we extended our symplectic point of view to the problem of constructing universal moduli spaces for Hitchin’s equations and Higgs bundles. The first steps of this attempt appeared in Chapter 7 of the fourth author’s 2018 ETH PhD Thesis [T], and is the seed of the present work. In view of the current interest in the field, after a dormant period, we decided to resume this investigation. We should warn that the use of the term ‘universal’ that we make here does not necessarily imply some expected functoriality properties that such term some times entails. We have sticked, however, to this term since that is the term that we originally used in our initial work [T].
After the completion of this work, we received a copy of a manuscript [CTW], where related topics are treated. It should be very interesting to compare our approach to other analytic constructions under study [CTW, Hi4], as well as the more algebraic point of view taken by Simpson [Si2, Si3] and others [BBN, BDP, D, DF].
@@ Acknowledgements.
The authors wish to thank Emilio Franco, Nigel Hitchin and Ignasi Mundet for helpful conversations. Special thanks are due to Richard Wentworth for stimulating discussions – in particular for spotting a mistake in the signature of the metric $\mathbf{g}^{\mathbb{I}}_{\alpha,1}$ in an earlier version of this manuscript – as well as for his lectures at ICMAT in June 2025, which encouraged us to resume this project.
@@ 2.1. Flat $\boldsymbol{G}$-connections
(2.2) $\Omega_{\mathbf{J}}(\dot{D}_{1},\dot{D}_{2})=\int_{X}B(\dot{D}_{1}\wedge\dot{D}_{2}),$
which is preserved by the action of the gauge group $\curly{G}$ of $E_{G}$. We have that $\operatorname{Lie}\curly{G}=\Omega^{0}(\Sigma,E_{G}(\mathfrak{g}))$ and hence its dual $(\operatorname{Lie}\curly{G})^{\ast}$ can be identified with $\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$. Similarly as for the case of unitary connections, studied by Atiyah–Bott [AB], there existes an equivariant complex moment map for the action of $\curly{G}$ on $\curly D$} given by
(2.3) $\mu{G}\colon{\curly D}$ $\longrightarrow$ $\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$
(2.4) $D$ $\longmapsto$ $F_{D},$
where $F_{D}$ is the curvature of $D$. The zero moment map condition, that is when $F_{D}=0$, corresponds to the flatness of $D$. In this case one refers to the pair $(E_{G},D)$ as a flat $G$-bundle.
In order to construct a Hausdorff moduli space of flat $G$-connections modulo gauge, one needs to impose a natural stability condition arising from Geometric Invariant Theory. We say that a bundle with $G$-connection $(E_{G},D)$ is reductive (also called completely reducible) if it splits as a product of irreducible sub-bundles, where irreducible (also called simple or stable in the literature) means that there exists no nontrivial $D$-invariant sub-bundle. For a flat $G$-bundle $(E_{G},D)$, this condition can be expressed in terms of the holonomy representation: the holonomy of $D$ defines a representation
$\rho:\pi_{1}(\Sigma)\to G,$
of the fundamental group $\pi_{1}(\Sigma)$ of $\Sigma$. A representation $\rho:\pi_{1}(X)\to G$ is said to be reductive if the Zariski closure of the image of $\rho$ is a reductive group, or equivalently if $\operatorname{ad}_{G}\circ\rho$ is completely reducible, where $\operatorname{ad}_{G}:G\to\mathrm{GL}(\mathfrak{g})$ is the adjoint representation. Then, one can prove that the flat connection $D$ on $E_{G}$ is reductive if and only if its corresponding representation is reductive.
@@
The moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ on $E_{G}$ admits a natural structure of singular complex manifold, and a holomorphic symplectic structure induced by $\Omega_{\mathbf{J}}$ on its smooth locus.
@@ 2.2. Surface group representations
$\pi_{1}(\Sigma)=\langle a_{1},b_{1},\dotsc,a_{g},b_{g}\;\;:\;\;\prod_{i=1}^{g}[a_{i},b_{i}]=1\rangle$
@@
Given a representation $\rho\colon\pi_{1}(\Sigma)\to G$, consider the associated flat $G$-bundle on $\Sigma$, defined as $E_{\rho}=\tilde{\Sigma}\times_{\rho}G$, where $\tilde{\Sigma}\to\Sigma$ is the universal cover and $\pi_{1}(\Sigma)$ acts on $G$ via $\rho$. This gives in fact an identification between the set of equivalence classes of representations $\operatorname{Hom}(\pi_{1}(\Sigma),G)/G$ and the set of equivalence classes of flat $G$-bundles, which in turn is parametrized by the cohomology set $H^{1}(X,G)$. We can then assign a topological invariant to a representation $\rho$ given by the characteristic class $c(\rho):=c(E_{\rho})\in\pi_{1}(G)$ corresponding to $E_{\rho}$. To define this, let $\widetilde{G}$ be the universal covering group of $G$. We have an exact sequence
@@
Let $E_{G}$ be a smooth principal $G$-bundle over $X$ with topological class $c$. Then there is a complex analytic isomorphism
@@
We warn the reader that, even though there is a complex analytic isomorphism between the moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ and the $G$-character variety with fixed topological invariant $c$, these spaces carry very different algebraic structures (see e.g. [Si3]).

### v2

@@ 1. Introduction
The moduli space $\mathcal{R}$ of representations of the fundamental group of a closed oriented surface $\Sigma$ has an interesting interplay with classical Teichmüller theory. For the case of a compact Lie group $K$, a choice of complex structure $J$ on $\Sigma$ determines a homeomorphism of the moduli space $\mathcal{R}(K)$ of $K$-representations with a moduli space of holomorphic principal $G$-bundles over the compact Riemann surface $X_{J}=(\Sigma,J)$, via the Narasimhan–Seshadri theorem [NS] and its extension to principal bundles by Ramanathan [Ra1, Ra2]. Here, $G$ denotes the complex reductive Lie group given by the complexification of $K$. This way, the character variety $\mathcal{R}(K)$ determines a holomorphic fibration over the Teichmüller space $\mathcal{T}$ of $\Sigma$, given by complex structures on $\Sigma$ modulo the action of diffeomorphisms isotopic to the identity.
@@
(1.1) $\begin{split}F_{A}-[\varphi,\tau(\varphi)]&=0,\\ \bar{\partial}_{J,A}\varphi&=0,\end{split}$
and subsequent work by Hitchin [Hi2, Hi3] and Simpson [Si1, Si2, Si3, Si4]. Here $A$ is a connection on a principal $K$-bundle over $\Sigma$ and $\varphi\in\Omega^{1,0}(X_{J},E_{K}(\mathfrak{g}))$ is the Higgs field. The corresponding gauge-theoretical moduli space $\mathcal{M}^{\operatorname{Hit}}(G)$ has three different incarnations, as a character variety $\mathcal{R}_{c}(G)$ for the complex reductive Lie group $G$, as a moduli space $\mathcal{M}^{\operatorname{Flat}}(G)$ of reductive flat $G$-connections, and as a moduli space of polystable $G$-Higgs bundles $\mathcal{M}^{Higgs}(G)$. These three viewpoints on the moduli space of Hitchin’s equations determine a hyperkähler structure on $\mathcal{M}^{\operatorname{Hit}}(G)$. Interestingly, only the moduli space $\mathcal{M}^{Higgs}(G)$ depends on the choice of a complex structure on $\Sigma$. Furthermore, the twistor space $Z\to\mathbb{CP}^{1}$ for the hyperkähler structure has generic complex structure $\mathcal{M}^{\operatorname{Flat}}(G)$, while the special points over $\mathbb{CP}^{1}$ correspond to $\mathcal{M}^{Higgs}(G)$ and its conjugate.
In this paper we investigate metric aspects of the universal moduli space of solutions of Hitchin’s equations varying over the Teichmüller space $\mathcal{T}$. Our approach is gauge-theoretical in nature, and builds on the classical theory of Kähler fibrations [GLS, Mu] and on the symplectic interpretation of constant scalar curvature Kähler metrics [Do2, Fu]. We fix a symplectic structure $\omega$ on $\Sigma$ and consider the coupled system of equations
@@
where $J$ is a complex structure on $\Sigma$ and $S_{g}$ is the scalar curvature of the metric $g=\omega(,J)$. Let $\mathcal{U}^{Hit}(G)$ be the moduli space of solutions of (1.2) modulo unitary gauge (see Section 5.4), and $\mathcal{U}^{Higgs}(G)$ be the universal moduli space of $G$-Higgs bundles (see Section 5.1). Our first main result can be stated as follows (cf. Theorem 5.18, and Theorem 5.22).
@@
(1.3) $\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mu_{\curly{H}}^{-1}(0)/\curly{H}$
@@
(1.5) $\begin{split}\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}\dot{J})\omega\\ &+\alpha\int_{\Sigma}B(a\wedge Ja)+\alpha\int_{\Sigma}B\left(\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\wedge J\left(\dot{\psi}+\tfrac{\alpha}{2}\psi(J\dot{J})\right)\right)\\ &-\frac{\alpha}{4}\int_{\Sigma}B\left(\psi(\dot{J})\wedge J\psi(\dot{J})\right),\end{split}$
where $B$ is a positive-definite invariant metric on $\mathfrak{k}=\operatorname{Lie}K$. Hence, it is not obvious a priori that $\mathcal{U}^{Hit}(G)$ inherits a complex structure compatible with $\boldsymbol{\omega}^{\mathbb{I}}_{\alpha,\varepsilon}$. The key step is to undertake a gauge fixing for solutions of the coupled Hitchin equations (1.2), whereby the complex structure $\mathbb{I}$ and the symmetric tensor $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ descend to the moduli space. Difficulties will arise, due to the fact that $\mathbf{g}^{\mathbb{I}}_{\alpha,\varepsilon}$ is neither a definite pairing nor non-degenerate.
@@
(1.7) $\mathcal{U}^{*}_{\alpha,\varepsilon}\longrightarrow\mathcal{U}^{\operatorname{Flat}}(G)$
@@
and $\nu{J}$ is induced by the global Kähler potential in the space of complex structures compatible with the orientation $\curly{J}$ (see [Fu, Section 4]).
Similarly to the universal moduli space of Hitchin’s equations (1.2), the (possibly degenerate) moduli space Kähler form and metric admit explicit formulae upon suitable gauge fixing
(1.8) $\begin{split}\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(J\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge\psi(\dot{J}_{2})),\\ \mathbf{g}^{\mathbb{J}}_{\alpha,\varepsilon}(v_{1},v_{2})&=\frac{\varepsilon}{2}\int_{\Sigma}\operatorname{tr}(\dot{J}_{1}\dot{J}_{2})\omega\\ &+\alpha\int_{\Sigma}B((a_{1}-\psi(\dot{J}_{1}))\wedge J(a_{2}-\psi(\dot{J}_{2})))\\ &+\alpha\int_{\Sigma}B((\dot{\psi}_{1}-(J\psi)(\dot{J}_{1}))\wedge J(\dot{\psi}_{2}-(J\psi)(\dot{J}_{2})))\\ &-\alpha\int_{\Sigma}B(\psi(\dot{J}_{1})\wedge J(\psi(\dot{J}_{2}))).\end{split}$
From the previous expression, we observe that, at least formally, the restriction of $\boldsymbol{\omega}^{\mathbb{J}}_{\alpha,\varepsilon}$ to the fibres over $\mathcal{T}$ coincides up to scaling with Hitchin’s symplectic structure $\omega_{\mathbf{J}}$. A complete proof of this requires a deeper understanding of the particular gauge fixing mechanism which we use to construct the moduli space complex structure $\mathbb{J}$ (see Proposition 4.9), and we leave it as an open question.
@@
We are far from even grasping the solution to this problem, and we leave it as an open question for future studies.
Motivation for the present work comes from the programme initiated by the first three authors in [AGG1] more than 15 years ago based on the 2009 PhD Thesis of the second author [Ga]. In that paper, we consider a pair $(X,E)$ consisting of a holomorphic principal $G$-bundle $E$ over a compact complex manifold $X$, where $G$ is a complex reductive group, and study certain coupled equations for a Kähler metric on $X$ and a reduction of structure group of $E$ to a maximal compact subgroup $K\subset G$. These equations, that we named Kähler–Yang–Mills equations, appear naturally as moment map equations for the action of the extended gauge group of the $K$-bundle $E_{K}$ over the underlaying smooth manifold to $X$ equipped with a symplectic form $\omega$. The extension of the gauge group of $E_{K}$ is given in this case by the group of Hamiltonian symplectomorphisms defined by $\omega$. This theory is in some sense a combination of the usual Yang–Mills theory and the Donaldson–Fujiki theory on Kähler manifolds, interpreting the constant scalar curvature as the moment map for the action of the group of Hamiltonian symplectomorphisms on the space of complex structures compatible with the symplectic structure $\omega$.
@@
In early joint work with the fourth author, we extended our symplectic point of view to the problem of constructing universal moduli spaces for Hitchin’s equations and Higgs bundles. The first steps of this attempt appeared in Chapter 7 of the fourth author’s 2018 ETH PhD Thesis [Tr], and is the seed of the present work. In view of the current interest in the subject, after a dormant period, we decided to resume this investigation. We should warn that the use of the term ‘universal’ that we make here does not necessarily imply any expected functoriality properties that such term some times entails. We have sticked, however, to this term since that is the one that we originally used in our initial work [Tr].
After the completion of this work, we received a copy of a manuscript [CTW], where related topics are treated. It should be very interesting to compare our approach to other analytic constructions under study [CTW, Hi4], as well as the more algebraic point of view taken by Simpson [Si2, Si3] and others [BBN, BDP, Da, DF].
@@ Acknowledgements.
The authors wish to thank Emilio Franco, Nigel Hitchin and Ignasi Mundet for helpful conversations. Special thanks are due to Richard Wentworth for stimulating discussions — in particular for spotting a mistake in the signature of the metric $\mathbf{g}^{\mathbb{I}}_{\alpha,1}$ in an earlier version of this manuscript — as well as for his lectures at ICMAT in June 2025, which encouraged us to resume this project.
@@ 2.1. Flat $\boldsymbol{G}$-connections
(2.2) $\Omega_{\mathbf{J}}(\dot{D}_{1},\dot{D}_{2})=\int_{\Sigma}B(\dot{D}_{1}\wedge\dot{D}_{2}),$
which is preserved by the action of the gauge group $\curly{G}$ of $E_{G}$. We have that $\operatorname{Lie}\curly{G}=\Omega^{0}(\Sigma,E_{G}(\mathfrak{g}))$ and hence its dual $(\operatorname{Lie}\curly{G})^{\ast}$ can be identified with $\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))$. As in the case of unitary connections, studied by Atiyah–Bott [AB], there exists an equivariant complex moment map for the action of $\curly{G}$ on $\curly D$} given by
(2.3) $\begin{array}[]{rcl}\mu{G}\colon{\curly D}&\longrightarrow&\Omega^{2}(\Sigma,E_{G}(\mathfrak{g}))\\ D&\longmapsto&F_{D},\end{array}$
where $F_{D}$ is the curvature of $D$. The zero moment map condition, that is, $F_{D}=0$, corresponds to the flatness of $D$. In this case one refers to the pair $(E_{G},D)$ as a flat $G$-bundle.
In order to construct a Hausdorff moduli space of flat $G$-connections modulo gauge, one needs to impose a natural stability condition arising from Geometric Invariant Theory. If the connection $D$ is flat, the holonomy map defines a representation $\rho$ of the fundamental group $\pi_{1}(\Sigma)$ of $\Sigma$ in $G$. The representation $\rho:\pi_{1}(\Sigma)\to G$ is said to be reductive if the Zariski closure of the image of $\rho$ is a reductive group, or equivalently if $\operatorname{ad}_{G}\circ\rho$ is completely reducible, where $\operatorname{ad}_{G}:G\to\mathrm{GL}(\mathfrak{g})$ is the adjoint representation. We say that $(E_{G},D)$ is reductive if $\rho$ is reductive.
@@
The moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ on $E_{G}$ admits a natural structure of a (possibly singular) complex manifold, and a holomorphic symplectic structure induced by $\Omega_{\mathbf{J}}$ on its smooth locus.
@@ 2.2. Surface group representations
$\pi_{1}(\Sigma)=\big\langle a_{1},b_{1},\dotsc,a_{g},b_{g}\;\;:\;\;\prod_{i=1}^{g}[a_{i},b_{i}]=1\big\rangle,$
@@
Given a representation $\rho\colon\pi_{1}(\Sigma)\to G$, consider the associated flat $G$-bundle on $\Sigma$, defined as $E_{\rho}=\tilde{\Sigma}\times_{\rho}G$, where $\tilde{\Sigma}\to\Sigma$ is the universal cover and $\pi_{1}(\Sigma)$ acts on $G$ via $\rho$. This gives in fact an identification between the set of equivalence classes of representations $\operatorname{Hom}(\pi_{1}(\Sigma),G)/G$ and the set of equivalence classes of flat $G$-bundles, which in turn is parametrized by the cohomology set $H^{1}(\Sigma,G)$. We can then assign a topological invariant to a representation $\rho$ given by the characteristic class $c(\rho):=c(E_{\rho})\in\pi_{1}(G)$ corresponding to $E_{\rho}$. To define this, let $\widetilde{G}$ be the universal covering group of $G$. We have an exact sequence
@@
Let $E_{G}$ be a smooth principal $G$-bundle over $\Sigma$ with topological class $c$. Then there is a complex analytic isomorphism
@@
We warn the reader that, even though there is a real analytic isomorphism between the moduli space of flat $G$-connections $\mathcal{M}^{\operatorname{Flat}}(G)$ and the $G$-character variety with fixed topological invariant $c$, these spaces carry very different complex structures (see e.g. [Si3]).